HIGH-RESOLUTION LASER SPECTROSCOPY OF A^{3}Π_{1u} ← X^{1}Σ_{g}^{+} SYSTEM OF Br_{2}

\underline{N.NISHIMIYA}, T.YUKIYA, and M.SUZUKI, Tokyo Polytechnic University, System Electronics and Information Technology, Iiyama 1583, Atsugi-City, 243-0297 Kanagawa, Japan.

Doppler limited vib-rotational spectrum of the A^{3}Π_{1u} ← X^{1}Σ_{g}^{+} system of ^{79}Br_{2}, ^{81}Br_{2}, and ^{79}Br^{81}Br have been measured using a titanium sapphire laser. We adapted the two-tone modulation method to detect the absorption spectrum^{1}. Measurements were made at room temperature in the region of 13000 – 13700 cm^{−1}and at 150^{∘}C in the region 12000 – 13000 cm^{−1}, as determined using a wavelength meter (Burleigh WA-1500) calibrated to the two-photon spectrum of Rb at 778 nm. The wavenumber accuracy measurement was estimated to be 0.005 cm^{−1}. About 10,000 lines of P, Q, and R-branches from 20 ≤ J ≤ 100 were assigned and spectroscopic constants of ^{79}Br^{81}Br for v′ = 4 − 9 were determined using a global least-squares fit to all lines of the three isotopic species by taking the reduced mass ratios into consideration. The details of the results will be presented.

1. D.E.Cooper and R.E.Warren, J.Opt.Soc.Am.B, 4, 470 (1987).